OpenStudy (help!!!!):

Help me get the steps for this word problem! The answer is 1:20pm but I keep on getting 1:15pm.

OpenStudy (help!!!!):

Carla and Juan need to paint 50 birdhouses. Carla can paint a birdhouse in 10 min. Juan can paint one in 6min. Carla gets started at 10am, but juan doesnt start until 10:20am. If they continue working until the job is complete, what time will they be done?

OpenStudy (bahrom7893):

How did u get 1:15 pm? Post your solution

OpenStudy (help!!!!):



OpenStudy (help!!!!):

@bahrom7893

OpenStudy (help!!!!):

@mathstudent55 @SolomonZelman @uri @ganeshie8 @ikram002p

OpenStudy (bahrom7893):

It looks right to me, I get 1:15 as well

OpenStudy (help!!!!):

Thanks for checking but the answer is really 1:20pm..

OpenStudy (bahrom7893):

If we start timing at 10:20 - carla had already painted 2 houses (10mins each), so: 10*t+6*t = 50-2 16t=48 t=48/16 = 6/2 = 3

OpenStudy (bahrom7893):

Wait no, how come the calculator showed 3.25, this is why calculators throw you off lol

OpenStudy (bahrom7893):

Yea, the answer's correct. 48/16 = (6*8)/(2*8) = 6/2 = 3hours

OpenStudy (bahrom7893):

Oh you probably added 2 instead of subtracting

OpenStudy (help!!!!):

hmm

OpenStudy (bahrom7893):

Yea, so you had 10x+6x-2 = 50 By which you mean, the number of houses painted by them both minus 2 should be 50

OpenStudy (bahrom7893):

We actually need the following: In total they had 50 houses to paint, but carla had already painted 2, leaving us with 50-2=48 remaining houses

OpenStudy (help!!!!):

OMG!! THank You!! The solved everything!

OpenStudy (bahrom7893):

So if we start counting when jose starts painting (at 10:20), they only have 48 houses left to paint 10x+6x=48 16x=48 x=3hours and if 3 hours passed after 10:20, you get 13:20, which is 1:20 pm

OpenStudy (bahrom7893):

cheers